Trace the curve y =³√ x −1 , and state all the properties you use to trace it.
1
Expert's answer
2020-03-04T17:35:52-0500

"y=f(x)=\sqrt[3]{x} - 1"

1) Domain

Domain is all real numbers.


2) Symmetry

"f(-x) = \sqrt[3]{-x}-1 = -\sqrt[3]{x}-1 \neq -f(x) \text{ or } f(x)"

The function is neither even nor odd.


3) Intercepts  on the axes

x-intercept:

"y=0 \\
\sqrt[3]{x} - 1 =0\\
x=1"

y-intercept:

"x=0\\
y=\sqrt[3]{0}-1=-1"


4) Asymptotes

Let "y=kx+b" is asymptote.

"k=\lim\limits_{x\to\infty}{\frac{f(x)}{x}}=\lim\limits_{x\to\infty}{\frac{\sqrt[3]{x}-1}{x}}=0"

"b=\lim\limits_{x\to\infty}{(f(x)-kx)}=\lim\limits_{x\to\infty}{\sqrt[3]{x}-1}=\infty"

No asymptotes.


5) Monotonicity

"y'=\frac{1}{3\sqrt[3]{x^2}}"

The first derivative does not exist when "x=0." If "x>0," then "y'>0." If "x<0," then "y'>0." Therefore, "y=f(x)" is monotonically increasing and "x=0" is the critical point.


6) Convexity

"y'' = -\frac{2}{9x\sqrt[3]{x^2}}"

The second derivative does not exist when "x=0." If "x>0," then "y''<0." If "x<0," then "y''>0." Therefore, when "x<0" the function is convex and when "x>0" the function is concave.